OpenStudy (anonymous):

Here is another question for you math scholars... -1/11+p=-7/3. Answer -74/33 How does this anwer compute?

myininaya (myininaya):

solve for x: a+x=b x=b-a solve for p: -1/11+p=-7/3 a+p=c p=c-a p=-7/3-(-1/11)=-7/3+1/11=(-77+3)/33=-74/33

OpenStudy (anonymous):

Will you write this out step by step and tell me how you came up with 33 being the denominator.

OpenStudy (anonymous):

Help me out here please?

myininaya (myininaya):

do you know how to find common denominator?